OpenStudy (please.help.me):

Find the exact values of sin(u/2), cos(u/2), and tan(u/2) using the half-angle formulas. cos u = 9/41, 0 < u < π/2 u/2 lies in quadrant 1

satellite73 (satellite73):

half angle formula here, you know it?

OpenStudy (please.help.me):

Yes!

satellite73 (satellite73):

\[\huge \sin(\frac{u}{2})=\sqrt{\frac{1-\cos(u)}{2}}\]

satellite73 (satellite73):

replace \(\cos(u)\) with \(\frac{9}{41}\)

OpenStudy (please.help.me):

So when I plug it in is u 9/41? I am confused on the plugging in

satellite73 (satellite73):

\[\sqrt{\frac{1-\cos(u)}{2}}\\ \sqrt{\frac{1-\frac{9}{41}}{2}}\]

satellite73 (satellite73):

hold up

satellite73 (satellite73):

it is not the cosine OF \(\frac{9}{41}\), the cosine IS \(\frac{9}{41}\)

OpenStudy (please.help.me):

What about the beginning sin(u/2) Will it be sin(9/41))/2 ?

satellite73 (satellite73):

ok i see this confusion frequently, so lets see if i can say it another way

satellite73 (satellite73):

you are told what the cosine of u is, i.e \[\cos(u)=\frac{9}{41}\] it is that number

satellite73 (satellite73):

all you do is replace \[\cos(u)\] in the formula by the number \(\frac{9}{41}\)

satellite73 (satellite73):

\[\huge \sin(\frac{u}{2})=\sqrt{\frac{1-\cos(u)}{2}}\] \[\color{red}{\cos(u)}=\color{red}{\frac{9}{41}}\] \[\huge \sin(\frac{u}{2})=\sqrt{\frac{1-\color{red}{\frac{9}{41}}}{2}}\]

OpenStudy (please.help.me):

I understand the problem and answer, but how would you write the beginning? cos(u/2) = ..... cos( <<<This part, like how do I plug that in?

OpenStudy (please.help.me):

Oh! Okay, so I guess I don't need to plug it in. It just stays the same?

satellite73 (satellite73):

you have three questions, sine, cosine and tangent of \(\frac{u}{2}\) i wrote the one for sine above

OpenStudy (please.help.me):

Yes! I understand the rest! Thank you!!!

satellite73 (satellite73):

i can try to say it differently but i don't know how to make it clear. the cosine of u, you are told that number you do not take the cosine of that number, it is the number itself

OpenStudy (please.help.me):

Wait!!!

OpenStudy (please.help.me):

Sin was actually marked wrong

satellite73 (satellite73):

maybe you have to do some more arithmetic

satellite73 (satellite73):

what did you write as your answer?

OpenStudy (please.help.me):

The very last screenshot you sent

satellite73 (satellite73):

that is only step one

OpenStudy (please.help.me):

This

satellite73 (satellite73):

you have to do the arithmetic

satellite73 (satellite73):

oh, also you put that is for cosine that should have been for sine

satellite73 (satellite73):

are you using webassign?

OpenStudy (please.help.me):

Yes!

satellite73 (satellite73):

it is very forgiving, maybe it will take that answer, but that answer was for sine, not for cosine you put it in the wrong box

satellite73 (satellite73):

you might also try \[\frac{4}{\sqrt{41}}\] which looks a lot better

OpenStudy (please.help.me):

Oh!!! Okay! Yes, it is marked correct for sine now

satellite73 (satellite73):

ok cosine is similar, but the minus is a plus

satellite73 (satellite73):

\[\huge \cos(\frac{u}{2})=\sqrt{\frac{1+\color{red}{\frac{9}{41}}}{2}}\]

OpenStudy (please.help.me):

Got it! Thank you so much!

satellite73 (satellite73):

you need tangent too right?

OpenStudy (please.help.me):

How do I find sin(x) for tan?

OpenStudy (please.help.me):

Reference triangle?

satellite73 (satellite73):

same way you always do draw a triangle

satellite73 (satellite73):



OpenStudy (please.help.me):

So 40?

satellite73 (satellite73):

there is actually an easier way, but yeah 40 i think

satellite73 (satellite73):

then \[\frac{1-\frac{9}{41}}{\frac{40}{41}}\]

OpenStudy (please.help.me):

sin(40) ________ 1+ (9/41)

satellite73 (satellite73):

grrr

satellite73 (satellite73):

it is not the SINE OF 40!!

satellite73 (satellite73):



OpenStudy (please.help.me):

Don't yell please

OpenStudy (please.help.me):

Thank you

satellite73 (satellite73):

\[\cos(u)=\frac{9}{41}\\ \sin(u)=\frac{40}{41}\]

satellite73 (satellite73):

do not take the sine or cosine of those numbers, those are the numbers you use ...(he said quietly)

OpenStudy (please.help.me):

Lol! Okay I got it! Thank you! Gave a great day!!! :)

satellite73 (satellite73):

your welcome, you too did you get it right?

OpenStudy (please.help.me):

@satellite73 Yes!

satellite73 (satellite73):

yay!!